OpenStudy (brandy):

We have an expense report with % of total and we have to find the angle to do a circle graph. The first is 25%, 20%, 8%, 5%,, 15%, 5%, 12%, 10%..and I am totally clueless on how to find the angle to do the circle graph.

OpenStudy (anonymous):

First convert each percent to a fraction 25% = 25/100 = 1/4 20% = 20/100 = 1/5 8% = 8/100 = 2/25 5% = 5/100 = 1/20 Now take each fraction and multiply them by 360º ( the number of degrees in a circle) 1/4 * 360º = 90º 1/5 * 360º = 72º 2/25 * 360º = 28.8º 1/20 * 360º = 18º and so on

OpenStudy (anonymous):

There are 360º in a circle. 25% of 360º is 0.25 x 360º = 90º. You can find the rest of the measurements that way as well. I always check to make sure the percentages add up to 100% as you would have trouble making a circle graph if they don't. You can also check that your angle measurements add up to 360º in the end. If they don't, you probably made a calculation error somewhere along the way.